hiiiihiiiiiiiiii:

Subtract this problem using mixed numbers. Reduce your answer to the lowest terms.12 4/5 - 4 1/25 16 5/30 9 3/5 8 3/25 8 19/25 I thank its C

hiiiihiiiiiiiiii:

I thank its C

hiiiihiiiiiiiiii:

beuase 12 - 4 = 8

ramen:

You have the right idea so far of the 8, but it's just the fraction part that is wrong.

hiiiihiiiiiiiiii:

ooooook

hiiiihiiiiiiiiii:

so then its D

ramen:

Yes and I'll explain. So you know how to find a common denominator right?

hiiiihiiiiiiiiii:

okkk

ramen:

So you were right on the part 12-4=8, but we are just left with the fractions 4/5 and 1/25. As you can see, the denominator 5 can go into 25 five times. So 5*5 is 25. Now we just multiply the 4*5 and we get 20. Instead of 4/5 we now have 20/25. understand so far?

hiiiihiiiiiiiiii:

yep

hiiiihiiiiiiiiii:

and also can you check this one

ramen:

Ok so all we have left is 20/25-1/25, 20-1= 19 and the denominator stays the same so we have 19/25. The answer is D: 8 19/25

hiiiihiiiiiiiiii:

. Subtract this problem using mixed numbers. Reduce your answer to the lowest term. 16 3/4 - 8 5/12 8 4/12 8 1/3 8 8 2/8 I thanks its D

ramen:

You can make a new post for that.

hiiiihiiiiiiiiii:

right